OpenStudy (anonymous):

If tan x ̊= 3/y and cos x ̊=y/z, what is the value of sin x°?

OpenStudy (anonymous):



OpenStudy (anonymous):

does that ugly picture help?

OpenStudy (anonymous):

the picture shows an angle \(x\) with \(\tan(x)=\frac{3}{y}\) and also \(\cos(x)=\frac{y}{z}\)

OpenStudy (anonymous):

use the ratio \(\sin(x)=\frac{\text{opposite}}{\text{hypotenuse}}\) to get your answer

OpenStudy (asnaseer):

an alternative approach is to note that:\[\tan(x)=\frac{\sin(x)}{\cos(x)}\]therefore:\[\sin(x)=\cos(x)\times\tan(x)\]

OpenStudy (anonymous):

that would certainly be snappier, wouldn't it?

OpenStudy (anonymous):

So is z/3 the answer?

OpenStudy (anonymous):

you have it upside down